OpenStudy (anonymous):

Find the equation of the horizontal line that passes through the point (-5,6).

OpenStudy (anonymous):

well if its horizontal that means that y is = to 0

OpenStudy (anonymous):

so the equation is y=0

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

what about this? Find the equation of the vertical line that passes through the point (-4,-3).

OpenStudy (anonymous):

then the x = 0

OpenStudy (anonymous):

y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

because when the line is verticle it touches the x axis

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

Find the equation of the line that has a slope of -3 and passes through the point (2,-2). how about this one?

OpenStudy (anonymous):



OpenStudy (anonymous):

well this one if you want to find the slope intercept form you can just say that m = slope b = y intercept so y = mx + b y = -3x + b then plug in the points -2 = -3(2) + b -2 = -6 + b 4 = b then you got your equation y = -3x + 4